The number of values of b for which there is an isosceles triangle with sides of length b + 5, 3b – 2 and 6 – b is
1. 0
2. 2
3. 1
4. 3

Grade:11

3 Answers

Ananya Verma
8 years ago
putting different values,
when b=2
property of isosceles triangle is proved.
kothapallisagarvivek
7 years ago
Case (I)
 b + 5 = 3b - 2
 => b = 7/2
 So sides are 5/2 , 17/2 , 17/2
 Case (II)
 b + 5 = 6 - b = b = 1/2
 Sides 11/2 , -1/2 , 11/2 is Not possible 
Case (III)
 3b - 2 = 6 - b 
4b = 8
 b = 2 
7, 4, 4 
two cases are possible
therefore answer is 2
kothapallisagarvivek
7 years ago
Case (I)
 b + 5 = 3b - 2
 => b = 7/2
 So sides are 5/2 , 17/2 , 17/2
 Case (II)
 b + 5 = 6 - b = b = 1/2
 Sides 11/2 , -1/2 , 11/2 is Not possible 
Case (III)
 3b - 2 = 6 - b 
4b = 8
 b = 2 
7, 4, 4 
two cases are possible
therefore answer is 2.